State of Residence        Frequency                      What proportion of this population lives

    Kansas                            5377                           in Nebraska?

    Missouri                            830

    Nebraska                          874

    Oklahoma                         500

    Colorado                          790

    Total                              8,371

        .104

        64.2

       10.4

       None of the above.

Question 2

    State of Residence        Frequency                      What percentage of this population lives

    Kansas                            5377                           in Missouri?

    Missouri                            830

    Nebraska                          874

    Oklahoma                         500

    Colorado                          790

    Total                               8371

                            

    .642

   .845

    9.9

    .059

Question 3

    State of Residence        Frequency                      What is the ratio of Oklahoma residents

    Kansas                            5377                           to Kansas residents?

    Missouri                            830

    Nebraska                          874

    Oklahoma                         500

    Colorado                          790

    Total                               8371

    10.754

    .092

    .059

    6.42
Answer #1

(1) proportion of this population lives in Nebraska = (Nebraska frequency)/(total frequency)

= (874/8371)

= 0.104

(2) proportion of this population lives in Missouri = [(Missouri frequency)/(total frequency)]*100

= (830/8371)*100

= 9.9

(3) ratio of Oklahoma residents  to Kansas residents = (Oklahoma frequency/Kansas frequency)

= (500/5377)

= 0.092